# Detect, Explain, Escalate: Low-Carbon Dialogue Breakdown Management for LLM-Powered Agents

### v2

@@ III-C Conversation-Level Labeling
In this case, the model seeks to predict after observing all $(u_{i},s_{i})$ pairs in $D$. This differs from utterance-level detection since any local breakdown event might cause the entire conversation to fail.
@@ IV Methodology
Our approach combines specialized fine-tuning and advanced prompting to achieve robust dialogue breakdown detection in both open-domain and task-oriented conversations. We focus on three key components: a compact fine-tuned model to detect dialogue breakdowns efficiently, leveraging advanced prompting strategies for generalist LLMs, and a proposed multi-tier inference architecture that calls upon frontier models only when necessary for corrections, optimizing both cost and accuracy.
@@ IV-A Supervised Fine-tuning with Reasoning Augmentation
To deploy an efficient alternative for real-time monitoring, we fine-tune a smaller model on labeled breakdown data using supervised fine-tuning (SFT). We choose Llama-3.1 8B [3] as it offers a good balance between accuracy and computational cost. The model is fine-tuned using the DBDC5 English and Japanese tracks [12] for per-utterance breakdown labels. Let:
  • $\mathcal{D}=\{(\mathcal{H}_{i},s_{i},b_{i})\}_{i=1}^{N}$ be the training data, where each sample has context $\mathcal{H}_{i}$, agent utterance $s_{i}$, and a binary label $b_{i}\in\{0,1\}$.
  • $T$ be a larger “teacher” LLM, such as Llama-3.3 70B that can generate synthetic reasoning traces $r_{i}$ (i.e., a chain-of-thought explanation) for each sample $(\mathcal{H}_{i},s_{i},b_{i})$.
We augment the original training input with this synthetic reasoning, effectively creating a more informative training set $\mathcal{D}^{\prime}$. We fine-tune our student model $S$ on the augmented dataset $\mathcal{D}^{\prime}$. The model is trained to predict the binary label $b_{i}$ (minimizing cross-entropy loss $\mathcal{L}_{CE}$) and to generate a textual justification $\hat{j}_{i}$. For the latter, the teacher-generated reasoning traces $r_{i}$ serve as training targets, enabling the student model to distill these structured reasoning patterns for improved classification and explanation. The cross-entropy loss [41] is defined as:
@@
where, $y_{i}$ denotes the true label for utterance $i$, and $\hat{y}_{i}$ denotes the predicted probability of breakdown. This training phase allows the student model to observe both the original dialogue context $(\mathcal{H}_{i},s_{i})$ and the teacher’s structured reasoning patterns $r_{i}$ during training, learning to distill these insights.
IV-B Advanced Prompting Strategies
While a fine-tuned compact model is efficient for real-time monitoring, powerful but costly LLMs can still be employed through strategic prompt engineering. Let $\mathcal{G}$ denote a generalist LLM, for instance GPT-4 or DeepSeek-R1. Given a conversation snippet $(\mathcal{H}_{i},s_{i})$, we form a prompt $\Pi(\mathcal{H}_{i},s_{i};\,\alpha)$ under a chosen strategy $\alpha\in\{\mathrm{ZS},\mathrm{FS},\mathrm{CoT},\mathrm{AR},\mathrm{CL+AR},\dots\}$. We then parse $g(\Pi(\mathcal{H}_{i},s_{i};\,\alpha))$ to obtain a breakdown label and confidence. The main prompting strategies we explore are:
Zero-Shot (ZS) Prompting. We provide only a task description and the current example $(\mathcal{H}_{i},s_{i})$. The model must infer the classification criterion from the instructions alone. A sample zero-shot prompt for DBDC5 is shown in Figure 1.
@@ IV-B Advanced Prompting Strategies
Few-Shot (FS) Prompting. We supply $k$ labeled examples of breakdown vs. non-breakdown before the new sample. Formally,
@@
Within FS, we explore:
  • 2-Shot Easy (2S-Easy): Two clear dialogues (one breakdown and one non-breakdown) where human annotators were very confident in the labels (>80% agreement). For example, an easy non-breakdown case might be a smooth dialogue that successfully completes, and an easy breakdown case might feature an obvious user hang-up after a system error. This helps the model anchor on unambiguous prototypes.
  • 2-Shot Hard (2S-Hard): We instead use two challenging samples where the correct outcome is subtle (annotator agreement 60–70%). One example might be a conversation with some confusion that eventually recovers (almost a breakdown, but not quite), and another might show a user hesitant or mildly frustrated (not a clear-cut hang-up, but dialogue quality is low). We hypothesize that exposing the model to ambiguous scenarios can improve its performance under uncertainty and lead to more calibrated confidence scores in similar situations.
  • 4-Shot (4S): We combine the Easy and Hard exemplars, two from 2S-Easy and two from 2S-Hard, thus exposing the model to a broader range of scenarios. While 4S can yield higher accuracy, prompt length grows, which increases token consumption and may approach context limits for some models, especially with longer dialogues.
    In datasets like BETOLD where per-utterance breakdown rationales from annotators are unavailable, we select short (15-20 turns) vs. long (21-30 turns) dialogues as “easier” vs. “harder” exemplars, then mimic the above FS setups accordingly. In addition, since the annotators did not provide their reasoning, we used Llama-3.3 70B to generate the step-by-step reasoning field given each exemplar, the annotators’ probability distribution (treated as a confidence score) and the decision label.
IV-C Self-Guided Reasoning Techniques
To further enhance the model’s reasoning depth and generalizability, we used and proposed advanced prompting techniques designed to encourage structured, self-generated reasoning:
Chain-of-Thought (CoT). We encourage multi-step reasoning by instructing $g$ to “think step-by-step” before deciding:
[…]
This often yields more coherent predictions on complex queries, albeit at higher token usage.
Analogical Reasoning (AR). Instead of providing examples directly, we ask $\mathcal{G}$ to generate hypothetical analogous conversations $\mathcal{A}_{i}$ from its internal knowledge, then classify the current conversation. Formally:
@@ IV-C Self-Guided Reasoning Techniques
where $\Pi_{\mathrm{AR}}$ instructs the model to “recall relevant past dialogues” similar to $(\mathcal{H}_{i},s_{i})$. The final prediction is then based on both $\mathcal{H}_{i},s_{i}$ and the newly generated analogies $\mathcal{A}_{i}$. This self-guided technique removes the need for hand-crafted exemplars, offering more tailored guidance for each test case.
IV-D Our Proposed Reasoning Technique
Curricular Learning with Analogical Reasoning (CL+AR). Curriculum learning involves training or prompting the model with an order of tasks that progresses from easy to hard, mimicking the way humans learn. The model builds a foundation to tackle more difficult ones by mastering simpler dialogue tasks. We propose applying this idea to LLM reasoning: the model is first steered to solve easy “proxy” queries related to the target query, then gradually given harder versions of the problem. The easy queries and solutions serve as stepping stones, forming a curriculum for the model’s chain of thought. In a dialogue context, we first ask the LLM to generate $\mathcal{A}_{i}^{(1)}$ (an easy analogous example), then $\mathcal{A}_{i}^{(2)}$ (moderate difficulty), and finally $\mathcal{A}_{i}^{(m)}$ (close to the given input’s complexity). This gradually “walks” the model from simpler to harder analogies. We then request the final decision for $(\mathcal{H}_{i},s_{i})$. Let $\mathcal{A}_{i}^{(1)},\dots,\mathcal{A}_{i}^{(m)}$ be the generated set of analogies. The overall prompt is:
@@ IV-D Our Proposed Reasoning Technique
We find that this method improves classification on challenging dialogues, although it may increase output token length.
IV-E Deployment Architecture
While a powerful LLM can reliably detect breakdowns, frequent usage may be cost-prohibitive. We address this via a real-time, hierarchical system (Figure 2) with three main modules to ensure dialogue robustness and maintain user trust:
@@ IV-E Deployment Architecture
    AI Assistant: Generates a candidate response $s_{i}$ given the user’s input $u_{i}$ and context $\mathcal{H}_{i}$:
    $s_{i}\;=\;\mathcal{G}_{\mathrm{assistant}}\!\bigl{(}\mathcal{H}_{i},\;u_{i}\bigr{)}.$
    Here, $\mathcal{G}_{\mathrm{assistant}}$ can be a moderately large language model such as Llama-3.3 70B.
@@
    Dialogue Disruption Monitor: Before presenting the assistant-generated response $s_{i}$ to the user, our fine-tuned model rapidly assesses the response for potential breakdowns or unsafe interactions, outputting:
@@
    where $\hat{b}_{i}\in\{0,1\}$ is the predicted breakdown label, $\hat{c}_{i}\in[0,1]$ is the confidence, and $\hat{j}_{i}$ (optional) is a justification. If $\hat{b}_{i}=0$ or $\hat{c}_{i}<T$, $s_{i}$ is deemed acceptable. Otherwise, we escalate to a larger “superior” LLM. The sub-second detection latency supports seamless user experiences.
@@
    Superior Model: Upon intervention, this model revises the response to prevent breakdown and ensure user safety:
    $s_{i}^{*}\;=\;\mathcal{G}_{\mathrm{superior}}\!\bigl{(}\mathcal{H}_{i},\;u_{i}\bigr{)}.$
    The monitor then re-checks $\bigl{(}\mathcal{H}_{i},\;s_{i}^{*}\bigr{)}$. This process can iterate up to $k$ times or until the breakdown is resolved.
This multi-tiered strategy significantly reduces computational demands and ensures resource-intensive models are only utilized in critical scenarios, thereby enhancing efficiency and user experience in conversational AI deployments.
@@ V Experiments
We evaluate our proposed approach using three dialogue breakdown detection datasets: English and Japanese tracks from DBDC5 and the BETOLD task-oriented dialogue dataset.
@@ V-A Dialogue Breakdown Datasets
Dialogue Breakdown Detection Challenge 5 (DBDC5). DBDC5, originating from the WOCHAT IWSDS 2020 workshop, serves as a widely recognized benchmark for dialogue breakdown detection research [12]. The evaluation set contains annotated dialogues in English (1,950 dialogues) and Japanese (2,672 dialogues), involving interactions between humans and conversational agents. Each dialogue turn is annotated by 15–30 human annotators. The system utterances are classified into three categories: Breakdown, Non-Breakdown, and Possible Breakdown (an intermediate category indicating uncertainty).
Breakdown Expectation for Task-Oriented Long Dialogues (BETOLD). Introduced in 2022, BETOLD identifies breakdowns specifically in task-oriented service dialogues, such as customer support calls [19]. This dataset comprises 13,524 real human–agent phone dialogues, annotated based on whether the interaction ended in a “late user-initiated hang-up or forward” (LUHF). LUHF instances indicate user frustration, leading to either terminating the call prematurely or requesting human intervention. Approximately 33% of the dataset represents breakdown cases. BETOLD dialogues utilize abstract representations of intents and entities derived from NLU/NLG annotations instead of raw textual data to maintain user privacy. We follow official train-test splits provided by dataset creators, treating BETOLD’s automatic labels as ground truth for evaluation purposes.
@@ V-B Data Preprocessing and Experimental Setup
Following prior works [9, 34], we preprocess dialogue data from DBDC5 and BETOLD datasets by converting annotations into binary labels (Breakdown/Non-Breakdown). Ambiguous annotations (‘Possible Breakdown’) are consolidated into the Breakdown class. To accommodate privacy constraints, the BETOLD dataset utilizes intent and entity abstractions rather than raw utterances, potentially limiting the efficacy of specific prompting techniques. We parse these structured dialogues and present them in a text form like “System: Intent: X — Entities: Y” for each turn, thereby preserving the sequence and dialogue flow in anonymized form. Basic cleaning (removal of extraneous symbols, ensuring consistent turn indexing) is applied for both datasets. No additional data augmentation is used beyond what is provided in the datasets.
@@
For advanced prompting methods (AR and CL+AR), we conducted experiments on a random subset representing 10% of each dataset due to higher token demands. Ablation studies indicated that a two-pass analogical reasoning approach (initial analogy generation followed by integration) performed comparably to single-pass prompting, aligning with findings in existing literature [25].
@@ V-C Fine-tuning Parameters
We fine-tune a Llama-3.1 8B instruct model using Low-Rank Adaptation (LoRA, rank=16) [42], employing an AdamW 8-bit optimizer [43] with a learning rate of $2\times 10^{-4}$, linear decay, batch size of 8, and weight decay of 0.01 across 3 epochs. Training utilized a single NVIDIA A100 40GB GPU. We hypothesize that a finely tuned smaller LLM can balance competitive accuracy with efficiency, making it viable for real-time dialogue monitoring. Our trained dialogue disruption monitor is openly available on HuggingFace [^1] https://huggingface.co/aghassel/dialogue_disruption_monitor, facilitating community engagement. Emphasizing sustainability, our fine-tuning approach, which leverages LoRA’s partial weight updates, significantly reduces computational load and energy consumption compared to traditional full-scale training.
@@ V-D Evaluation Metrics
We use accuracy and F1 score as our primary evaluation metrics for breakdown detection. Accuracy measures the overall correctness of breakdown vs non-breakdown predictions. F1 score (the harmonic mean of precision and recall) is especially important because the breakdown class is under‑represented; it evaluates how well models detect breakdowns without neglecting the non‑breakdown class. We therefore report the F1 score for both classes along with overall accuracy, in line with prior work on DBDC5 [32, 34]. This mirrors the official challenge metrics and provides continuity with previous state-of-the-art results. Metrics are computed on held-out test sets (DBDC5 evaluation sets and reserved BETOLD splits). We parse the model’s JSON-formatted output to extract the decision label. In cases where model outputs deviate from the required JSON format, instead of discarding these responses, we employ Llama-3.3 70B as an LLM-based judge to interpret the model’s textual output and extract the intended classification decision when the JSON format is not strictly adhered to.
On the DBDC English and Japanese tracks, we investigated the model’s overconfidence using the MSE between the annotators’ probability distribution for a dialogue utterance and the LLM’s verbalized confidence in its corresponding decision.

### v3

@@ III-C Conversation-Level Labelling
Here the model predicts a single label after observing all $(u_{i},s_{i})$ pairs in $D$. Conversation-level breakdown often results from one or more local breakdown events, but the relationship is not necessarily one-to-one.
@@ IV Methodology
Our approach combines supervised fine-tuning with advanced prompting strategies to achieve accurate dialogue breakdown detection in both open-domain and task-oriented settings. The framework comprises three components: (1) a compact fine-tuned model (Llama-3.1 8B) for efficient real-time breakdown detection, (2) prompting strategies (few-shot, chain-of-thought, and analogical reasoning) for large-scale LLMs, and (3) a multi-tier inference architecture that escalates to high-capacity models only when confidence thresholds indicate potential breakdowns.
@@ IV-A Supervised Fine-tuning with Reasoning Augmentation
To support real-time monitoring, we fine-tune a smaller model on labelled breakdown data using supervised fine-tuning (SFT). We select Llama-3.1 8B [llama] as a balance between accuracy and computational cost. The model is fine-tuned on the DBDC5 English and Japanese tracks [higashinaka_dialogue_2016] for per-utterance breakdown labels.
Let:
  • $\mathcal{D}=\{(\mathcal{H}_{i},s_{i},b_{i})\}_{i=1}^{N}$ denote the training data, where each sample comprises a context $\mathcal{H}_{i}$, agent utterance $s_{i}$, and a binary label $b_{i}\in\{0,1\}$.
  • $T$ be a larger teacher LLM, such as Llama-3.3 70B, which generates synthetic reasoning traces $r_{i}$ (chain-of-thought style explanations) for each sample $(\mathcal{H}_{i},s_{i},b_{i})$.
We augment the training inputs with these synthetic reasoning traces, forming an enriched dataset $\mathcal{D}^{\prime}$. The student model $S$ is fine-tuned on $\mathcal{D}^{\prime}$ to (i) predict the binary breakdown label $b_{i}$ by minimizing cross-entropy loss and (ii) generate a textual justification $\hat{j}_{i}$. Teacher-generated traces $r_{i}$ serve as targets for justification, encouraging the student to internalize structured reasoning patterns that improve both classification and interpretability. The cross-entropy loss [shannon_cel] is:
@@
where $y_{i}$ is the true breakdown label and $\hat{y}_{i}$ is the predicted breakdown probability. During training, the student observes both the original dialogue context $(\mathcal{H}_{i},s_{i})$ and the teacher’s reasoning $r_{i}$, enabling distillation of richer decision rules than supervision on labels alone.
IV-B Prompting Strategies
While a fine-tuned compact model is suitable for continuous monitoring, larger LLMs can still be leveraged for higher-fidelity assessment under carefully designed prompts. Let $\mathcal{G}$ denote a generalist LLM such as GPT-4 or DeepSeek-R1. Given $(\mathcal{H}_{i},s_{i})$, we form a prompt $\Pi(\mathcal{H}_{i},s_{i};\,\alpha)$ under strategy $\alpha\in\{\mathrm{ZS},\mathrm{FS},\mathrm{CoT},\mathrm{AR},\mathrm{CL+AR},\dots\}$. The model output is parsed into a breakdown label and confidence.
Zero-Shot (ZS) Prompting. In the zero-shot setting, we provide a natural language task description and the current example $(\mathcal{H}_{i},s_{i})$. The model must infer the decision boundary from instructions alone. A sample DBDC5 zero-shot prompt is shown in Figure 1.
@@ IV-B Prompting Strategies
Chain-of-Thought (CoT). CoT prompts encourage multi-step reasoning by instructing $\mathcal{G}$ to “think step by step” before deciding:
$\Pi_{\mathrm{CoT}}(\mathcal{H}_{i},s_{i})\;=\;\bigl[\mathcal{H}_{i},\,s_{i},\,\text{``Let’s think step by step.''}\bigr].$
This often improves predictions on complex cases, though it increases token usage.
Few-Shot (FS) Prompting. Few-shot prompts prepend $k$ labelled exemplars before the test example:
@@
We consider several variants:
  • 2-Shot Easy (2S-Easy): two clear dialogues (one breakdown, one non-breakdown) with high annotator agreement ($80\%$). For example, an easy non-breakdown case might be a smooth dialogue that successfully completes, and an easy breakdown case might feature an obvious user hang-up after a system error. This helps the model anchor on unambiguous prototypes.
  • 2-Shot Hard (2S-Hard): exemplars with moderate annotator agreement (60–70%), representing borderline cases. These can improve calibration by exposing the model to uncertainty. One example might be a conversation with some confusion that eventually recovers (almost a breakdown, but not quite), and another might show a user mildly frustrated (not a clear-cut hang-up, but dialogue quality is low).
  • 4-Shot (4S): a mix of two easy and two hard exemplars, providing broader coverage at the cost of longer prompts.
For BETOLD, which lacks per-utterance rationales, we select dialogues by length (15–20 vs. 21–30 turns) as proxies for easier and harder cases. Since the annotators did not provide their reasoning, we generate the step-by-step reasoning field for each example using Llama-3.3 70B, given the annotators’ probability distribution (treated as a confidence score) and the decision label.
Analogical Reasoning (AR). In analogical prompting, we instruct $\mathcal{G}$ to construct hypothetical dialogues that are analogous to $(\mathcal{H}_{i},s_{i})$ and then classify the original example. Formally,
@@
where $\Pi_{\mathrm{AR}}$ asks the model to “recall or construct past dialogues similar to this one.” The final decision uses both the original dialogue and generated analogies $\mathcal{A}_{i}$. This self-generated context removes the need for manually crafted exemplars.
Curriculum Learning with Analogical Reasoning (CL+AR). Curriculum learning organizes examples from easy to hard. We apply this idea to analogical prompting by asking the LLM to generate a sequence of analogies $\mathcal{A}_{i}^{(1)},\dots,\mathcal{A}_{i}^{(m)}$ that gradually increase in difficulty, culminating in the target-like scenario:
@@
The hope is that solving simpler analogous cases first helps the model reason more reliably about the true example.
IV-C Deployment Architecture
While large-scale LLMs demonstrate reliable breakdown detection capabilities, their computational cost and inference latency make continuous deployment impractical for real-time applications. We therefore adopt a hierarchical architecture (Figure 2) with three modules:
@@ IV-C Deployment Architecture
    AI Assistant: A base assistant model generates a candidate response $s_{i}$ given the user’s input $u_{i}$ and history $\mathcal{H}_{i}$:
    $s_{i}\;=\;\mathcal{G}_{\mathrm{assistant}}\!\bigl(\mathcal{H}_{i},\;u_{i}\bigr),$
    where $\mathcal{G}_{\mathrm{assistant}}$ may be a moderately sized LLM such as Llama-3.3 70B.
@@
    Dialogue Disruption Monitor: Before returning $s_{i}$ to the user, our fine-tuned monitor evaluates it for potential breakdowns or unsafe behaviour:
@@
    where $\hat{b}_{i}$ is the breakdown label, $\hat{c}_{i}$ is a confidence score, and $\hat{j}_{i}$ is an optional justification. If $\hat{b}_{i}=0$ and $\hat{c}_{i}<T$, we accept $s_{i}$. Otherwise, we escalate.
@@
    Superior Model: When escalation is triggered, a more capable model revises the response:
    $s_{i}^{*}:=\mathcal{G}_{\mathrm{superior}}\!\bigl(\mathcal{H}_{i},u_{i},\hat{j}_{i}\bigr),$
    producing a corrected response $s_{i}^{*}$ using the history, user input, and monitor’s justification. Conditioning on $\hat{j}_{i}$ (e.g., “response is contradictory”) allows the superior model to apply targeted repairs rather than recomputing from scratch.
In customer-service scenarios, a high-confidence breakdown prediction ($\hat{c}_{i}>T$) could additionally trigger a silent alert to a human agent. The alert includes $\hat{j}_{i}$, enabling efficient human review and intervention. While the superior model could in principle act as both assistant and monitor, doing so at every turn would negate the efficiency gains of the hierarchy. The lightweight monitor is thus essential for scalable and sustainable deployment.
@@ V Experiments
We evaluate our approach on three dialogue breakdown detection benchmarks: the English and Japanese tracks of DBDC5 and the BETOLD dataset for task-oriented dialogues.
@@ V-A Dialogue Breakdown Datasets
Dialogue Breakdown Detection Challenge 5 (DBDC5). DBDC5, introduced at the WOCHAT IWSDS 2020 workshop, is a standard benchmark for dialogue breakdown detection [Higashinaka2020DBDC5, higashinaka_dialogue_2016]. The evaluation set includes 1,950 English and 2,672 Japanese dialogues between humans and conversational agents. Each system utterance is annotated by 15–30 human annotators into three categories: Breakdown, Non-Breakdown, and Possible Breakdown. The agents span rule-based, retrieval-based, and neural generative architectures. Although these differ from modern LLM-based systems, the underlying breakdown phenomena (topic drift, contradiction, ignored context) remain salient sources of error [Huang2025Taxonomy].
Breakdown Expectation for Task-Oriented Long Dialogues (BETOLD). Introduced in 2022, BETOLD targets breakdowns in real-world, task-oriented customer-service calls [terragni2022_BETOLD]. It contains 13,524 human–agent phone dialogues, annotated according to whether the interaction ends with a “late user-initiated hang-up or forward” (LUHF), indicating user frustration and either call termination or escalation. Around 33% of dialogues are labelled as breakdowns. To protect privacy, BETOLD encodes dialogues as sequences of intents and entities rather than raw text. We follow the official train–test splits and treat the provided labels as ground truth.
@@ V-B Data Preprocessing and Experimental Setup
We binarize labels on DBDC5 and BETOLD, merging Possible Breakdown with Breakdown, following [Ghassel2024Are, lin_semi-supervised_2022]. To accommodate privacy constraints, the BETOLD dataset utilizes intent and entity abstractions rather than raw utterances, potentially limiting the efficacy of specific prompting techniques. We parse these structured dialogues and present them in a text form like “System: Intent: X — Entities: Y” for each turn, thereby preserving the sequence and dialogue flow in anonymized form. Basic cleaning (removal of extraneous symbols, ensuring consistent turn indexing) is applied for both datasets. No additional data augmentation is used beyond what is provided in the datasets.
@@
For LLM-based experiments we require models to output three fields: a justification, a decision, and a confidence score. We found that the order of these outputs affects behaviour: prompting for a justification first and a decision second encourages more deliberative reasoning and leads to better calibration than requesting the label upfront. For AR and CL+AR, we restrict evaluation to 10% subsamples of each dataset due to high token cost. Preliminary experiments indicate that a two-pass analogical pipeline (analogy generation followed by decision) performs comparably to single-pass prompting, aligning with existing literature [analogical].
@@ V-C Fine-tuning Parameters
We fine-tune Llama-3.1 8B using Low-Rank Adaptation (LoRA, rank = 16) [hu2022lora]. We use AdamW with 8-bit parameters [adamw], a learning rate of $2\times 10^{-4}$, linear decay, batch size 8, weight decay 0.01, and train for three epochs on a single NVIDIA A100 40GB GPU. We hypothesize that a carefully fine-tuned small LLM can provide competitive accuracy while remaining suitable for real-time monitoring. Our trained dialogue disruption monitor is released on HuggingFace [^1] https://huggingface.co/aghassel/dialogue_disruption_monitor. Because LoRA updates only a subset of parameters, this training regime substantially reduces compute and energy consumption compared to full-parameter fine-tuning.
@@ V-D Evaluation Metrics
We report accuracy and F1 scores for both Breakdown (B) and Non-Breakdown (NB) classes. Accuracy measures overall correctness; class-specific F1 scores capture performance on the minority Breakdown class without ignoring the majority NB class. Following DBDC5 challenge practice [ng_improving_2020, lin_semi-supervised_2022], we report accuracy and per-class F1 on held-out evaluation sets (DBDC5 official evaluation and reserved BETOLD test splits).
Models are prompted to output JSON-formatted responses. We parse these outputs to extract the decision label. When outputs deviate from the expected format, we employ Llama-3.3 70B as an LLM-based judge to interpret the textual response and recover the intended classification.
On DBDC5 English and Japanese, we also assess overconfidence by computing MSE between the model’s verbalized confidence and the annotator-derived breakdown probabilities.